OpenStudy (anonymous):

\[cot{{-5\pi}\over6}\]

OpenStudy (anonymous):

-sqrt3?

OpenStudy (anonymous):

sqrt3*

OpenStudy (anonymous):



OpenStudy (anonymous):

cot(−5π/6)=-cot(5π/6)=-cot(π -π/6)=-(-cotπ/6)=+cotπ/6=sqrt(3)

OpenStudy (anonymous):

thanks! :)